189-235A: Basic Algebra I Assignment 1

Due: Wedday, September 11

Appendix B: 9, 14, 20, 27, 28, 33. Appendix C: 9, 13.

9. Evaluate the complex numbers (1 + i)^{64} and (1 + \sqrt{3}i)^{72}.

10. Prove that ∑_{k=0}^{n} (\frac{n}{k}) = 2^{n}, and that ∑_{k=0}^{n}(−1)^{k} (\frac{n}{k}) = 0, using the binomial theorem.